# From Simulations to Reality: Dark Energy Reconstruction with Simulated SNIa data from the Vera C. Rubin Observatory

### v2

@@ Abstract
In this paper, we present an Artificial Neural Network (ANN) based reconstruction analysis of the Supernova Ia (SNIa) distance moduli ($\mu(z)$), and hence dark energy, using LSST simulated three-year SNIa data. Our ANN reconstruction architecture can model both the distance moduli and their corresponding error estimates. For this we employ astroANN and incorporate Monte Carlo dropout techniques to quantify uncertainties in our predictions. We tune our hyperparameters through advanced genetic algorithms, including elitism, utilizing the DEAP library. We compared the performance of the ANN based reconstruction with two theoretical descriptions of dark energy models, $\Lambda$CDM and Chevallier-Linder-Polarski (CPL). We perform a Bayesian analysis for these two theoretical models using the LSST simulations and also compare with observations from Pantheon and Pantheon+ SNIa real data. We show that our model-independent reconstruction using ANN is consistent with both of them. We assessed the performance using mean squared error (MSE) and showed that the ANN can produce distance estimates in better agreement with the LSST dataset than either $\Lambda$CDM or CPL, albeit very small. We included an additional residual analysis and a null test with $F$-scores to show that the reconstructed distances from the ANN model, are in excellent agreement with the $\Lambda$CDM or CPL model.
@@ I Introduction
The discovery of the accelerated expansion of our Universe using Supernova Ia observations revolutionized our understanding of modern cosmology (riess_1998, ; Perlmutter:1998np, ; 2008ApJ…686..749K, ; Betoule2014, ; Planck2016, ; 2012ApJ…746…85S, ). The standard model of cosmology, $\Lambda$CDM, is composed of ordinary matter, a dark energy modeled as a cosmological constant $\Lambda$ that is responsible for the acceleration of the Universe, and Cold Dark Matter (CDM) Carroll:2000fy that shapes the cosmic structure through gravitational influence. $\Lambda$CDM assumes a homogeneous and isotropic Universe and has been in excellent agreement with most of the currently available data. However, it has its theoretical drawbacks such as fine-tuning and cosmic coincidence, and from an observational point of view, it also suffers from the Hubble tension (DiValentino:2021izs, ). These issues open the door to study models beyond the standard cosmological model.
@@
There are non-parametric inference techniques that try to avoid assumptions of specific theoretical models, to find new dark energy properties from the data. They extract information directly from the data and infer unknown quantities based mainly on the data with as few assumptions as possible sahni2002cosmological ; wasserman2006all . Some interesting examples of these types of methods previously used to reconstruct cosmological functions are Principal Component Analysissharma2020reconstruction , smoothed step functions r8 , Gaussian processes williams2006gaussian ; Keeley:2020aym ; l2020defying ; r10 ; escamilla2023model and extrapolation methods montiel2014nonparametric . These approaches are also known as model-independent reconstructions, and they can be considered as new models based on the data, they can be used to analyze their similarity with different theoretical models and, to determine what model describes better the observational data. There are several works based on model-independent approaches to analyze dark energy features sahni2006reconstructing ; holsclaw2010nonparametric ; zhao2017dynamical ; r8 , cosmic expansion montiel2014nonparametric , deceleration parameter r10 , the growth rate of structure formation l2020defying ; gomez2023neuralrecs and luminosity distance wei2017improved ; lin2019non ; wang2020reconstructing .
@@
In current times, the increase in computing power, bigger telescopes, enhanced CCDs, and a large amount of observational data have allowed the incorporation of machine learning methods as analysis tools in observational cosmology lin2017does ; peel2019distinguishing ; arjona2020can ; wang2020machine ; gomez2023neuralgenetic ; chacon2023analysis ; medel2023cosmological . Artificial Neural Networks (ANN) are one of the most revolutionary methods of Artificial Intelligence (AI), and they have been successfully employed in a wide range of applications in cosmology, for example, in image analysis dieleman2015rotation ; ntampaka2019deep , N-body simulations chacon2023analysis ; rodriguez2018fast ; he2019learning and statistical methods gomez2023neuralrecs ; gomez2023neuralgenetic ; alsing2019fast ; li2019model ; hortua2020constraining ; hortua2020parameter . From the data perspective, the forthcoming observational programs are anticipated to enhance the dark energy precision by almost ten folds FoM , some examples are Euclid EuclidTheoryWorkingGroup:2012gxx ; https://doi.org/10.48550/arxiv.0912.0914 , LSST LSSTDarkEnergyScience:2018jkl ; Zhan:2017uwu ; LSST:2008ijt , Roman Space Telescope Akeson:2019biv , the Thirty Meter Telescope (TMT) TMT , and the already operational JWST Gardner_2006 . Consequently, all the reasons above culminate in the increasing effort of leveraging AI and large datasets to investigate dark energy in a new way. This approach aims to overcome the longstanding issues of model degeneracy and the tension between different cosmological measurements without relying on predetermined theoretical models.
@@
In this paper, we present a dark energy reconstruction analysis based on the state-of-the-art LSST’s three years of simulated PLAsTiCC SNIa data plasticc_announce and Artificial Neural Network modeling. We generated neural network models for the distance modulus and we compared this model with the Bayesian parameter estimation of dark energy models. In addition, we contrast the results with the SNIa data from the Pantheon and Pantheon+ surveys as real observational data. Our Bayesian analysis allows us to explore whether a model preference can be inferred from the LSST dataset, between the two most prevalent formulations of dark energy: $\Lambda$CDM and CPL (see §II.1). This paper is structured as follows, in sec. II we present the theory of the two different dark energy models we adopted in this analysis and a brief overview of the ANNs. In sec. III we outline the LSST survey program, which leads us to the next section IV, where we introduce the LSST simulated SNIa dataset; which is the primary dataset for our analysis. We also introduce the additional Pantheon SNIa datasets later in this section, used for comparisons only (and no ANN training is done with them). In the next sec. V, we discuss the methods used for constructing the ANN network and how the model is trained. In sec. VI we present the findings of our analysis and finally sec. VII presents the conclusion of this paper.
@@ II.1 Dark Energy Models
and a cosmological constant $\Lambda$ corresponds to $w=-1$ and $\rho$ (energy density) constant. Over the last two decades, the $\Lambda$CDM model has consistently been shown to have good agreement with the observations. The Hubble parameter relates to the redshift ($z$) as,
@@
it can be seen as a first-order approximation of the Taylor expansion of a more general dark energy EoS $w(a)$ (2017ApJ…850..183Z, ). This approach ensures that the model remains well-behaved across the entire range of the Universe’s history, from high to low redshifts. The dynamics of the Universe’s expansion are encapsulated by the scale factor $a$, with the background evolution described as,
@@
The CPL is very well-behaved, all the way, from $a=0$ to $a=1$. Being a simple two-parameter model, the CPL is easy to analyze and also it adapts well to a wide variety of dark energy models.
@@ II.2 Supernova Ia
Type Ia supernovae (SN Ia) are believed to be the result of the thermonuclear disruption of a carbon-oxygen white dwarfs which reaches the Chandrasekhar-mass limit of stability ($M_{Ch}\sim 1.4$M solar mass) by accreting matter from a companion 1960ApJ…132..565H . SNIa light curves show remarkable homogeneity after correcting for stretch and color parameters. Observed flux from SNIa can be used to compute the luminosity distance $D_{L}$. For an FLRW Universe, with a dark energy component EoS $w(z)$ the $D_{L}$ is given as,
@@
and the distance modulus $\mu$ can be related to the $D_{L}$ (for a flat Universe with a constant EoS dark energy $w=-1$) as
$\mu=5\log(D_{L}/10\text{pc}),$ (7)
therefore SNIa provides a direct and robust method to probe dark energy. Over the last two decades, numerous studies have utilized Type Ia Supernovae (SNIa) as the primary probe for investigating dark energy panstarrs ; Betoule2014 ; s1 ; Dhawan:2021ztf . Through successive SNIa programs, our precision in understanding dark energy has significantly improved, with efforts still ongoing Kessler2010 ; Linder2019 ; Mitra2021 .
@@ II.3 Dark Energy Reconstruction and Artificial Neural Networks
While parametric reconstruction methods have provided valuable insights by imposing specific models on the dark energy EoS $w(z)$, they inherently assume a certain level of theoretical bias toward the functional form of dark energy as discussed above. These limitations fanned the study of non-parametric reconstruction methods, which offer a more model-independent avenue for analyzing dark energy. Unlike their parametric counterparts, non-parametric methods do not rely on predefined equations or functions to describe the reconstructed parameter r8 ; r9 ; r10 ; r11 ; r12 . Instead, they leverage the data directly, allowing the underlying properties to manifest more freely in the analysis. In context to dark energy, this shift towards non-parametric reconstruction approaches can significantly open up new possibilities by reducing the theoretical prejudices and using the complexity and richness of the observational data r13 ; r14 ; gomez2023neuralgenetic . In this analysis, we use ANN-based non-parametric reconstruction technique gomez2023neuralrecs ; Wang:2020sxl ; Wang:2019vxv to obtain distance modulus ($\mu$) from SNIa observations. ANNs are powerful computational models capable of approximating any continuous nonlinear function, as demonstrated by hornik1990universal , making them exceptionally suited for modeling complex and large datasets. They are also better than the Gaussian Process (GP) based reconstructions in the sense that the ANN approach offers enhanced flexibility (eg. overfitting issues, kernel selection sensitivity, etc.) and accuracy in modeling natural processes compared to GP by having more optimizable hyperparameters. They are also less influenced by any prior assumptions which GP is more sensitive of dialektopoulos2022neural . For a comprehensive background on neural networks, seminal texts such as goodfellow2016deep ; bishop2006pattern ; nielsen2015neural provide in-depth analyses. The details of our ANN-based architecture are outlined in sec.§V.
@@ III LSST: Overview
The Legacy Survey of Space and Time (LSST), conducted by the Vera C. Rubin Observatory collaboration, represents a paradigm shift in astrophysical surveys with its unparalleled scope and technical sophistication. Scheduled to begin in 2023, the LSST will deploy an advanced observational apparatus, featuring an $8.4$ m primary mirror with a $6.7$ m effective aperture and a state-of-the-art $3200$-megapixel camera, yielding a wide field of view of $9.6$ square degrees. The LSST is designed to survey approximately 18,000 square degrees of the southern sky over a decade, utilizing six optical passband filters to facilitate deep, wide, and fast observations. The survey aims to amass over 32 trillion observations of 20 billion galaxies and a similar number of stars, achieving a depth of $24^{th}$ magnitude in its six filter bands, spanning wavelengths from ultraviolet to near-infrared Cahn2009 ; ivezic . These efforts are projected to catalog millions of supernovae, among other transient phenomena, offering an unprecedented dataset for probing the dark Universe. The Dark Energy Science Collaboration (DESC)  [^2] https://lsstdesc.org/, comprising nearly 1,000 members, intends to harness this vast trove of data to extract high-precision measurements of fundamental cosmological parameters, leveraging prior data challenges to refine analysis pipelines in anticipation of the survey’s extensive data output (dc2, ; Sanchez, ).
@@ IV Datasets
Our reconstruction analysis using the ANN models is based on the LSST 3-year simulated SNIa data. We have, also for comparison, analyzed the performance of the considered dark energy models using Pantheon and Pantheon+ SNIa datasets, all of which are described in this section below.
@@ IV.1 Simulated LSST SNIa Data
The SN data is generated using the LSST DESC time domain (TD) pipeline and SNANA code (snana, ), consisting of four main stages (illustrated in Fig. 2(a) of mitra2023using ). These include SN brightness standardization via a Light Curve (LC) fit stage, simulations for bias correction, and a BBC stage for Hubble diagram production before the last stage, cosmology fitting. We used a Hubble diagram and the associated covariance matrix (statistical + systematic) produced from the BBC stage to perform the cosmological fitting.
@@
For mock generation, input cosmology: $\Omega_{m}=0.3150,\,\Omega_{\Lambda}=0.6850,\,w_{0}=-1,\,w_{a}=0$ was used. The curvature is computed internally as $\Omega_{k}=1-\Omega_{m}-\Omega_{\Lambda}$. It is based on cosmological parameters from Planck 2018 Planck2018 . In addition, the parameter $H_{0}$ is set to $70.0$ km/s/Mpc, this value is tied to SALT2 training salt and we use the SALT2 lightcurve model. It generates observer frame magnitudes. The noise in the simulation is computed as follows :
$\sigma_{\text{SIM}}^{2}=\left[F+(A\cdot b)+(F\cdot\sigma_{\text{ZPT}})^{2}\right.\left.+\sigma_{0}\cdot 10^{0.4\cdot ZPT_{\text{pe}}}+\sigma_{\text{host}}^{2}\right]S^{2}_{\text{SNR}},$ (8)
where $F$ is the simulated flux in photoelectron (p.e.), $A$ is the noise equivalent area given by, $A=\left[2\pi\int\text{PSF}^{2}(r,\theta)r\mathop{}\!\mathrm{d}r\right]^{-1}$ where PSF stands for the Point Spread Function, $b$ is the background per unit area (includes sky + CCD readouts + dark current), $S_{\text{SNR}}$ is an empirically determined scale that depends on the signal-to-noise ratio, the three $\sigma$ terms correspond to zero point uncertainty, flux calibration uncertainty and underlying host galaxy uncertainty. These terms can be determined empirically from fits that match simulated uncertainties to those from the survey designs.
@@
We use a redshift-binned Hubble diagram (HD) alongside its corresponding covariance matrix. The datasets are composed of a mixture of spectroscopically ($z_{\rm spec}$) and photometrically ($z_{\rm phot}$) identified SNIa candidates at low redshift and high redshift respectively.
@@
The photometric redshift determination and its uncertainty were based on Graham2018_photoz while using host galaxy photo-$z$ as priors (Kessler2010 ). mitra2023using re-simulated the PLAsTiCC  data based on the DDF strategy of the LSST, augmented with low-redshift spectroscopic data from the DC2 analysis. The study by Mitra:2022ykq , which simulated exclusively SNIa without considering contamination, generated a statistical plus systematic covariance matrix to account for seven systematics. The HD is comprised of 5809 SNIa candidates. Importantly, the work by Mitra:2022ykq expanded on the initial analysis plan for the LSST outlined by LSSTDarkEnergyScience:2018jkl , which focused on a 1 and 10-year timeline for studying Supernova Type Ia (SNIa) cosmology. Mitra:2022ykq did this by conducting a comprehensive analysis that takes into account both the spectroscopic and photometric redshifts of the supernova’s host galaxies.

### v3

# Dark energy reconstruction analysis with artificial neural networks: Application on simulated Supernova Ia data from Rubin Observatory

@@ Abstract
In this paper, we present an analysis of Supernova Ia (SNIa) distance moduli ($\mu(z)$) and dark energy using an Artificial Neural Network (ANN) reconstruction based on LSST simulated three-year SNIa data. The ANNs employed in this study utilize genetic algorithms for hyperparameter tuning and Monte Carlo Dropout for predictions. Our ANN reconstruction architecture is capable of modeling both the distance moduli and their associated statistical errors given redshift values. We compare the performance of the ANN-based reconstruction with two theoretical dark energy models: $\Lambda$CDM and Chevallier-Linder-Polarski (CPL). Bayesian analysis is conducted for these theoretical models using the LSST simulations and compared with observations from Pantheon and Pantheon+ SNIa real data. We demonstrate that our model-independent ANN reconstruction is consistent with both theoretical models. Performance metrics and statistical tests reveal that the ANN produces distance modulus estimates that align well with the LSST dataset and exhibit only minor discrepancies with $\Lambda$CDM and CPL.
@@ I Introduction
The discovery of the accelerated expansion of our Universe using Supernova Ia observations revolutionized our understanding of modern cosmology (riess_1998, ; Perlmutter:1998np, ; 2008ApJ…686..749K, ; Betoule2014, ; Planck2016, ; 2012ApJ…746…85S, ). The standard model of cosmology, $\Lambda$CDM, is composed of ordinary matter, a dark energy modeled as a cosmological constant $\Lambda$ that is responsible for the acceleration of the Universe, and Cold Dark Matter (CDM) Carroll:2000fy that shapes the cosmic structure through gravitational influence. $\Lambda$CDM assumes a homogeneous and isotropic Universe and has been in excellent agreement with most of the currently available data. However, it has its theoretical drawbacks, such as fine-tuning and cosmic coincidence SolaPeracaula:2022hpd , and observationally it suffers from the Hubble tensions between the values of $H_{0}$ obtained from CMB observations Planck2018 and SNIa observations pantheon_new ; Pantheon ; Riess_2022 ; DiValentino:2021izs ; Aluri:2022hzs ; Perivolaropoulos:2021jda ; Kamionkowski:2022pkx ; DiValentino:2021izs ; Krishnan:2021dyb ; Abdalla:2022yfr ; Krishnan:2021jmh ; Smith:2022hwi . These issues open the door to study models beyond the standard cosmological model.
@@
There are non-parametric inference techniques that try to avoid assumptions of specific theoretical models, to find new dark energy properties from the data. They extract information directly from the data and infer unknown quantities based mainly on the data with as few assumptions as possible sahni2002cosmological ; wasserman2006all . Some interesting examples of these types of methods previously used to reconstruct cosmological functions are Principal Component Analysissharma2020reconstruction , smoothed step functions r8 , Gaussian processes williams2006gaussian ; Keeley:2020aym ; l2020defying ; r10 ; escamilla2023model and extrapolation methods montiel2014nonparametric . These approaches are also known as model-independent reconstructions, and they can be considered as new models based on the data, they can be used to analyze their similarity with different theoretical models and, to determine what model describes better the observational data. There are several works based on model-independent approaches to analyze dark energy features sahni2006reconstructing ; holsclaw2010nonparametric ; zhao2017dynamical ; r8 , cosmic expansion montiel2014nonparametric , deceleration parameter r10 , the growth rate of structure formation l2020defying ; gomez2023neuralrecs and luminosity distance wei2017improved ; lin2019non ; wang2020reconstructing . In particular, Dark Energy has been studied using neural networks gomez2023neuralrecs ; escamilla2020deep ; wang2020reconstructing ; dialektopoulos2023neural , genetic algorithms arjona2020machine ; arjona2020can ; arjona2020hints ; arjona2022testing , and other machine learning techniques escamilla2023model ; Keeley:2020aym .
@@
In current times, the increase in computing power, bigger telescopes, enhanced CCDs, and a large amount of observational data have allowed the incorporation of machine learning methods as analysis tools in observational cosmology lin2017does ; peel2019distinguishing ; arjona2020can ; wang2020machine ; gomez2023neuralgenetic ; chacon2023analysis ; medel2023cosmological . Artificial Neural Networks (ANN) are one of the most revolutionary methods of Artificial Intelligence (AI), and they have been successfully employed in a wide range of applications in cosmology, for example, in image analysis dieleman2015rotation ; ntampaka2019deep , N-body simulations chacon2023analysis ; rodriguez2018fast ; he2019learning and statistical methods gomez2023neuralrecs ; gomez2023neuralgenetic ; alsing2019fast ; li2019model ; hortua2020constraining ; hortua2020parameter .
We are using neural networks to generate model-independent models based on the data with a methodology presented in previous works gomez2023neuralrecs ; gomez2023neuralgenetic . A key advantage of well-trained neural networks is their independence from a fiducial cosmology, allowing the generated data to be treated as new observations reflecting the true nature of the original dataset. Furthermore, neural networks surpass standard interpolation techniques by not requiring any statistical distribution assumptions for the data due to their nonlinear modeling capabilities. From the data perspective, the forthcoming observational programs are anticipated to enhance the dark energy precision by almost ten folds FoM , some examples are Euclid EuclidTheoryWorkingGroup:2012gxx ; https://doi.org/10.48550/arxiv.0912.0914 , LSST LSSTDarkEnergyScience:2018jkl ; Zhan:2017uwu ; LSST:2008ijt , Roman Space Telescope Akeson:2019biv , the Thirty Meter Telescope (TMT) TMT , and the already operational JWST Gardner_2006 . Consequently, all the reasons above culminate in the increasing effort of leveraging AI and large datasets to investigate dark energy in a new way. This approach aims to overcome the longstanding issues of model degeneracy and the tension between different cosmological measurements without relying on predetermined theoretical models.
@@
In this paper, we present a dark energy reconstruction analysis based on the state-of-the-art LSST’s three years of simulated PLAsTiCC SNIa data plasticc_announce and Artificial Neural Network modeling. We generated data-based models with neural networks for the distance modulus and we compared this model with the Bayesian parameter estimation of dark energy models. In addition, we contrast the results with the SNIa data from the Pantheon and Pantheon+ surveys as real observational data. Our Bayesian analysis allows us to explore whether a model preference can be inferred from the LSST dataset, between the two most prevalent formulations of dark energy: $\Lambda$CDM and CPL (see §II.1).
This paper is structured as follows, in sec. II we present the theory of the two different dark energy models we adopted in this analysis and a brief overview of the ANNs. In sec. III we outline the LSST survey program, which leads us to the next section IV, where we introduce the LSST simulated SNIa dataset; which is the primary dataset for our analysis. We also introduce the additional Pantheon SNIa datasets later in this section, used for comparisons only (and no ANN training is done with them). In the next sec. V, we discuss the methods used for constructing the ANN network and how the model is trained. In sec. VI we present the findings of our analysis and finally sec. VII presents the conclusion of this paper.
@@ II.1 Dark Energy Models
and a cosmological constant $\Lambda$ corresponds to $w=-1$ and $\rho$ (energy density) constant. Over the last two decades, the $\Lambda$CDM model has consistently been shown to have good agreement with the observations. For a flat Universe, the Hubble parameter relates to the redshift ($z$) as,
@@
it can be seen as a first-order approximation of the Taylor expansion of a more general dark energy EoS $w(a)$ (2017ApJ…850..183Z, ). This approach ensures that the model remains well-behaved across the entire range of the Universe’s history, from high to low redshifts. The dynamics of the Universe’s expansion are encapsulated by the scale factor $a$, with the background evolution described as (for a flat Universe),
@@
The CPL is very well-behaved, all the way, from $a=0$ to $a=1$. Being a simple two-parameter phenomenological description of dark energy, the CPL is easy to analyze and also it adapts well to a wide variety of dark energy models.
@@ II.2 Supernova Ia
Type Ia supernovae (SN Ia) are believed to be the result of the thermonuclear disruption of carbon-oxygen white dwarfs which reaches the Chandrasekhar-mass limit of stability ($M_{Ch}\sim 1.4$M solar mass) by accreting matter from a companion 1960ApJ…132..565H . SNIa light curves show remarkable homogeneity after correcting for stretch ($x_{1}$, i.e. the time stretching of the light curve) and SN color ($c$) at the peak brightness parameters. Observed flux from SNIa can be used to compute the luminosity distance $D_{L}$. For an FLRW Universe, with a dark energy component EoS $w(z)$ the $D_{L}$ is given as,
@@
Distance estimations in SNIa are based on an empirical observations, that all these events belong to a homogeneous class, whose variablilities can be captured mostly via the stretch and the color component. For each event, the measured distance modulus ($\mu$) is based on the Tripp formula Trip1998 ,
$\mu=m_{B}+\alpha x_{1}-\beta c+M_{0}+\Delta\mu_{\rm bias}~{},$ (7)
where $m_{B}$ ($=-2.5\log_{10}(x_{0})$) is the approximate B-band peak brightness magnitude of the SN. $m_{B}$ along with $x_{1}$ and $c$ are the SALT3 light curve parameters. While $\alpha$ and $\beta$ are global nuisance parameters which sets the amplitude of the stretch-luminosity and color-luminosity corrections respectively. $M_{0}$ is the absolute magnitude of the SNIa, with $x_{1}=0$, $c=0$, and is also a global nuisance parameter. Without a calibrated absolute distance scale, $M_{0}$ is degenerate with the Hubble parameter $H_{0}$. The final term $\Delta\mu_{\rm bias}=(\mu-\mu_{true})$, accounts for the corrections resulting from selection effects and analysis biases and is determined from a separate biasCor simulation in a 5 dimensional space of $\{z,x_{1},c,\alpha,\beta\}$ bbc .
Over the last two decades, numerous studies have utilized Type Ia Supernovae (SNIa) as the primary probe for investigating dark energy panstarrs ; Betoule2014 ; s1 ; Dhawan:2021ztf . Through successive SNIa programs, our precision in understanding dark energy has significantly improved, with efforts still ongoing Kessler2010 ; Linder2019 ; Mitra2021 .
@@ II.3 Dark Energy Reconstruction and Artificial Neural Networks
While parametric reconstruction methods have provided valuable insights by imposing specific models on the dark energy EoS $w(z)$, they inherently assume a certain level of theoretical bias toward the functional form of dark energy as discussed above. These limitations fanned the study of non-parametric reconstruction methods, which offer a more model-independent avenue for analyzing dark energy. Unlike their parametric counterparts, non-parametric methods do not rely on predefined equations or functions to describe the reconstructed parameter r8 ; r9 ; r10 ; r11 ; r12 . Instead, they leverage the data directly, allowing the underlying properties to manifest more freely in the analysis. In context to dark energy, this shift towards non-parametric reconstruction approaches can significantly open up new possibilities by reducing the theoretical prejudices and using the complexity and richness of the observational data r13 ; r14 ; gomez2023neuralgenetic . In this analysis, we use ANN-based non-parametric reconstruction technique gomez2023neuralrecs ; Wang:2020sxl ; Wang:2019vxv to obtain distance modulus ($\mu$) from SNIa observations. ANNs are powerful computational models capable of approximating any continuous nonlinear function, as demonstrated by hornik1990universal , making them exceptionally suited for modeling complex and large datasets. They are also better than the Gaussian Process (GP) based reconstructions in the sense that the ANN approach offers enhanced flexibility (eg. overfitting issues, kernel selection sensitivity, etc.) and accuracy in modeling natural processes compared to GP by having more optimizable hyperparameters. They are also less influenced by any prior assumptions which GP is more sensitive of dialektopoulos2022neural .
For a comprehensive background on neural networks, seminal texts such as goodfellow2016deep ; bishop2006pattern ; nielsen2015neural provide in-depth analyses. The details of our ANN-based architecture are outlined in sec.§V.
@@ III LSST: Overview
The Legacy Survey of Space and Time (LSST), conducted by the Vera C. Rubin Observatory collaboration, represents a paradigm shift in astrophysical surveys with its unparalleled scope and technical sophistication. Scheduled to begin in 2025, the LSST will deploy an advanced observational apparatus, featuring an $8.4$ m primary mirror with a $6.7$ m effective aperture and a state-of-the-art $3200$-megapixel camera, yielding a wide field of view of $9.6$ square degrees. The LSST is designed to survey approximately 18,000 square degrees of the southern sky over a decade, utilizing six optical passband filters to facilitate deep, wide, and fast observations. The survey aims to amass over 32 trillion observations of 20 billion galaxies and a similar number of stars, achieving a depth of $24^{th}$ magnitude in its six filter bands, spanning wavelengths from ultraviolet to near-infrared Cahn2009 ; ivezic . These efforts are projected to catalog millions of supernovae, among other transient phenomena, offering an unprecedented dataset for probing the dark Universe. The Dark Energy Science Collaboration (DESC)  [^2] https://lsstdesc.org/, comprising nearly 1,000 members, intends to harness this vast trove of data to extract high-precision measurements of fundamental cosmological parameters, leveraging prior data challenges to refine analysis pipelines in anticipation of the survey’s extensive data output (dc2, ; Sanchez, ).
@@ IV Datasets
Our reconstruction analysis using the ANN models is based on the LSST like a 3-year simulated SNIa data. We have, also for comparison, analyzed the performance of the considered dark energy models using Pantheon and Pantheon+ SNIa datasets, all of which are described in this section below.
@@ IV.1 Simulated LSST SNIa Data
The SN data is generated using the LSST DESC time domain (TD) pipeline and SNANA code (snana, ), consisting of four main stages (illustrated in Fig. 4 of mitra2023using ). These include SN brightness standardization via a Light Curve (LC) fit stage, simulations for bias correction, and a BBC stage for Hubble diagram production before the last stage, cosmology fitting. We used a Hubble diagram and the associated covariance matrix (statistical + systematic) produced from the BBC stage to perform the cosmological fitting.
@@
For mock generation, input cosmology: $\Omega_{m}=0.3150,\,\Omega_{\Lambda}=0.6850,\,w_{0}=-1,\,w_{a}=0$ was used. The curvature is computed internally as $\Omega_{k}=1-\Omega_{m}-\Omega_{\Lambda}$ (for our analysis case this is consistent with the assumptions of a flat cosmology). It is based on cosmological parameters from Planck 2018 Planck2018 . In addition, the parameter $H_{0}$ is set to $70.0$ km/s/Mpc, this value is tied to SALT2 training salt and we use the SALT2 lightcurve model [^3] Like mentioned in sec. II.2 that $H_{0}$ has a degeneracy with $M_{0}$ and hence the choice of $H_{0}$ value is not significant..
@@
We use a Hubble diagram (HD) containing redshift, distance modulus information alongside its corresponding covariance matrix. The datasets are composed of a mixture of spectroscopically ($z_{\rm spec}$) and photometrically ($z_{\rm phot}$) identified SNIa candidates at low redshift and high redshift respectively.
@@
The photometric redshift determination and its uncertainty were based on Graham2018_photoz while using host galaxy photo-$z$ as priors (Kessler2010 ). mitra2023using re-simulated the PLAsTiCC  data based on the DDF strategy of the LSST  [^4] https://www.lsst.org/scientists/survey-design/ddf, augmented with low-redshift spectroscopic data from the DC2 analysis Sanchez . Based on PLAsTiCC  plasticc_announce $z_{\rm spec}$” sample is composed of two sets of events whose spectroscopic redshifts have an accuracy of $\sigma_{z}\sim 10^{-5}$. The first subset is made up of spectroscopically confirmed events whose redshift is predicted to be accurate by the 4MOST spectrograph  4MOST2 , which is being built by the European Southern Observatory and is expected to become operational in $2023$. 4MOST is situated at a latitude similar to that of the Rubin Observatory in Chile. The second subset is composed of photometrically identified events with an accurate host galaxy redshift determined by 4MOST. The second subset has about $\sim 60$% more events than the first subset. For the photometric sample, host galaxy photo-$z$ was used as a prior (adapted from  Kessler2010 ). The photometric redshift and rms uncertainty was based on  Graham2018_photoz . The whole simulation was re-done based on the PLAsTiCC  DDF data [^5] The LSST has different observing strategies, the deep field and the wide field, called as the DDF (Deep drilling field) and the WFD (Wide Fast Deep) respectively.. Additional low redshift spectroscopic data was used from the DC2 analysis (simulated with WFD cadence  [^6] While WFD cadence is being used for the DC2 analysis, a full scale WFD study is yet not done, and is underway currently in the LSST.). The study by mitra2023using , which simulated exclusively SNIa without considering contamination, generated a statistical plus systematic covariance matrix to account for seven systematics. The HD is comprised of 5785 SNIa candidates. Importantly, the work by mitra2023using expanded on the initial analysis plan for the LSST outlined by LSSTDarkEnergyScience:2018jkl , which focused on a 1 and 10-year timeline for studying Supernova Type Ia (SNIa) cosmology. Since this study is based on three years simulated SNIa from the LSST using the DDF observations only hence the smaller sample size, as opposed to millions that is forecast to observe during its ten years study using both the DDF and the Wide-Fast-Deep (WFD) observations combined. mitra2023using did this by conducting a comprehensive analysis that takes into account both the spectroscopic and photometric redshifts of the supernova’s host galaxies.